OpenStudy (anonymous):

The area of a rectangle is 28 m2 , and the length of the rectangle is 1 m more than double the width. Find the dimensions of the rectangle.

OpenStudy (abhisar):

Let the breadth be x then length will be (2x+1). Area of rectangle = length x Breadth => x(2x+1) = 28 Solve for x and find the length and breadth

OpenStudy (anonymous):

i tried doing that but i kept on getting the answer wrong

OpenStudy (abhisar):

Can you show me your work so that I could correct you

OpenStudy (anonymous):

sure how do i upload a picture

OpenStudy (abhisar):

Use the attach file below

OpenStudy (anonymous):

it is not letting me

OpenStudy (anonymous):

nevermore i'll just get this one wrong

OpenStudy (abhisar):

Wait, You will get an equation \(\sf 2x^2 +x-28 = 0\)

OpenStudy (abhisar):

Now, you need to solve this quadratic equation. You will get two values for x, one positive and one negative.

OpenStudy (anonymous):

so i solve for x

OpenStudy (abhisar):

Yes, You can use this website to check your value of x http://www.math.com/students/calculators/source/quadratic.htm

OpenStudy (abhisar):

Note:- You need to pick the positive value. The value of x will be your breadth and the length will be 2x+1

OpenStudy (anonymous):

ok I think i understand I'm going to try to do it :)

OpenStudy (abhisar):

Ok c: